Boadu and Ansah formed a company and agreed that their annual profit will be shared in the ratio 4:5 respectively. If at the end of the year, Ansah received GHC 5,000.00 more than Boadu. How much was Boadu's share?
Answer by ikleyn(52787)

.
Boadu and Ansah formed a company and agreed that their annual profit will be shared in the ratio 4:5 respectively.
If at the end of the year, Ansah received GHC 5,000.00 more than Boadu. How much was Boadu's share?
~~~~~~~~~~~~~~~~

Their shares are  4x (Boadu)  and  5x (Ansah),  where x is the common factor.


We are given that  5x - 4x = 5000.

Therefore,  x = 5000.


It implies that Boadu's share is 4x = 4*5000 = 20000.    ANSWER

Solved.